# Reinforcement Learning with Verifiable Rewards: GRPO’s Effective Loss, Dynamics, and Success Amplification

### v3

@@ 2.3. Stabilized GRPO with No Clipping
We will focus in what follows on this non-clipped version.
@@
Algorithm 1 Iterative GRPO with verifiable rewards, modified from (Shao et al., 2024)
@@
8:    Compute $A(q,o_{i})$ using equation (3), where $p=p_{\theta_{\text{old}}}(q)=\frac{1}{G}\sum_{i=1}^{G}\mathbbm{1}_{r(q,o_{i})=1}$
@@
10:         Update the policy model $\pi_{\theta}$ by maximizing (GRPO-No-Clip)) with gradient ascent (Equation
11:Output $\pi_{\theta}$
2.4. GRPO Iterations
Algorithm 1 summarizes GRPO iterations (Stabilized and no clipping). We see that GRPO iterations can be written as a sequence optimization we denote by $\pi_{\theta_{n}}$, the policy at iteration $n$. We see that GRPO iterations can be written for $n\geq 1$:
Note that in Algorithm 1, expectations are estimated using importance sampling from $\pi_{\theta_{n-1}}$, and each maximization problem is solved via gradient for $\mu$ steps.
$p_{n-1}(q)=\mathbb{E}_{o\sim\pi_{n-1}(.|q)}\mathbbm{1}_{r(q,o)=1}$ (7)
and the weights $\omega^{+}_{\varepsilon}$ and $\omega^{-}_{\varepsilon}$ are given in Equation (4). We assume all throughout the paper that $\pi_{0}=\pi_{\mathrm{ref}}.$ Note that moving the optimization from a parametric space to the probability space can be seen as assuming that the hypothesis class of the parametric policies is large enough to represent all policies. Note that in GRPO iterations the policy at iteration $n$ depends upon the policy $\pi_{n-1}$ via the probability of success $p_{n-1}$, as well on the reference policy via the $\mathsf{KL}$ regularizer.
3. GRPO Dynamics: Fixed Point iteration for Probability of Success
Our goal in this Section is to analyze the dynamics of the GRPO iterations given in Equation (GRPO Iterations).
Theorem 1 (GRPO Policy Dynamic).
Optimal GRPO iterations policies solving Equation (GRPO Iterations) satisfy the following recursion, for $n\geq 1$:
where
$Z_{n-1}(q)=p_{\mathrm{ref}}(q)\exp\left(\frac{1}{\beta}\omega^{+}_{\varepsilon}(p_{n-1}(q))\right)+(1-p_{\mathrm{ref}}(q))\exp\left(-\frac{1}{\beta}\omega^{-}_{\varepsilon}(p_{n-1}(q))\right),$
where the weights $\omega^{+}_{\varepsilon}$ and $\omega^{-}_{\varepsilon}$ are given in Equation (4), the probability of success $p_{n-1}(q)$ of policy $\pi_{n-1}(\cdot|q)$ is given in Equation (7), and $p_{\mathrm{ref}}(q)$ is the probability of success of the reference policy $\pi_{\mathrm{ref}}(\cdot|q)$: $p_{\mathrm{ref}}(q)=\mathbb{E}_{o\sim\pi_{\mathrm{ref}}(\cdot|q)}\mathbbm{1}_{r(q,o)=1}$.
Theorem 2 (GRPO’s Probability of Success Fixed Point Iteration).
Assume $p_{\mathrm{ref}}>0$, define for $\beta>0$:
$p_{n}(q)=h_{\varepsilon,p_{\mathrm{ref}}}(p_{n-1}(q)),$ (8)
and $p_{0}(q)=p_{\mathrm{ref}}(q)$.
Remark 1 (Importance of $\varepsilon>0$).
Note if $\varepsilon=0$, $h_{\varepsilon,p_{\mathrm{ref}}}$ is no longer continuous at 00 and $1$ and we can no longer guarantee existence of fixed points on $[0,1]$.
We study in the following proposition propreties of the function $h_{\varepsilon,p_{\mathrm{ref}}}$:
Proposition 1 (Propreties of $h_{\varepsilon,p_{\mathrm{ref}}}$).
$h_{\varepsilon,p_{\mathrm{ref}}}$ satisfies the following propreties:
  • Existence of fixed points: $h_{\varepsilon,p_{\mathrm{ref}}}$ is continuous $[0,1]$ to $[0,1]$ and hence admits at least a fixed point $p^{*}$ in $[0,1]$ (no guarantees for a unique fixed point)
We drop in the sequel $q$, when referring to the sequence $p_{n}(q)$, and write for short $p_{n}$ (the reader is referred to Remark 2 for a discussion). If the sequence defined in GRPO’s probability of success iteration (8) converges we have therefore by continuity of $h_{\varepsilon,p_{\mathrm{ref}}}$:
4. GRPO: Fixed Point Iteration Convergence and Success Amplification
In this Section we answer the following two questions:
  1. (1)
    Do fixed points of GRPO iterations, $p^{*}$, lead to a probability of success $p^{*}$ that is higher than the reference initialization $p_{0}=p_{\mathrm{ref}}$?
  2. (2)
    Under which conditions do we have local convergence of the GRPO’s probability of success sequence given in (8) to a fixed point $p^{*}$ of $h_{\varepsilon,p_{\mathrm{ref}}}$?
Theorem 3 (GRPO amplifies the probability of success).
Let $10<p_{\mathrm{ref}}<$ 1. Let $p^{*}$ be a fixed point of $h_{\varepsilon,p_{\mathrm{ref}}}$ we have $p^{*}>p_{\mathrm{ref}}$.
We see from Theorem 3 that the fixed point $p^{*}$ of the GRPO iteration leads to an amplification of the probability of success of the reference model $p_{\mathrm{ref}}$. Note if $p_{\mathrm{ref}}=0$ or $p_{\mathrm{ref}}=1$ , the iteration will lead to $p^{*}=0$ and $p^{*}=1$ respectively.
We now turn to the second question regarding the convergence of the GRPO sequence given in (8) to a fixed point $p^{*}$ of $h_{\varepsilon,p_{\mathrm{ref}}}$. Given the properties of $h_{\varepsilon,p_{\mathrm{ref}}}$, we can characterize the limit point of the GRPO iteration as $n\to\infty$ as follows, as a consequence of the local Banach fixed-point theorem:
Theorem 4 (Local Fixed Point Convergence).
$\lim_{n\to\infty}p_{n}=p^{*}.$
Lemma 1.
$=p^{*}(1-p^{*})\frac{2p^{*}-1}{2\beta\,[p^{*}(1-p^{*})+\varepsilon]^{3/2}}$
We see from Figure 3 the lower bound on $\beta$ required to ensure local convergence of GRPO iterations to a fixed point $p^{*}$. Figure 4 shows iteration (8) as a function of $n$ for different values of $\beta$ and $p_{\mathrm{ref}}$. We see that in most cases, there is a sharp transition where we observe fast convergence to $1$ or to a fixed point $p^{*}$. For $\beta=5$ and $p_{\mathrm{ref}}=0.001$, we see a divergent behavior.
Remark 2.
5. Back to Parametric GRPO Iterations
Assumption 1.
We have the following theorem:
Theorem 5.
$\lim_{n\to\infty}|\tilde{p}_{n}-p^{*}|\leq 2\delta^{*}.$
In the case $\delta^{*}=0$, we have convergence to the fixed point.
6. Experimental Validation
Setup
We use the GSM8K dataset from Cobbe et al. (2021) (MIT license), and Qwen2.5-0.5B-Instruct (Apache 2.0 license) by Yang et al. (2024) as the reference policy. We use GRPO implememtation in TRL (von Werra et al., 2020b), and train on the training split of GSM8K on a node with 8 GPUs (GPU0 for the vLLM server and 7 other GPUs for distributed training). We use a learning $5e^{-6}$ and the KL regularizer $\beta=0.1$, and $\mu$ in Algorithm 1 is set to $\mu=10$. Other hyperparameters are given in Appendix B . We use the correctness of the LLM output as a reward.
Success Rate Amplification
The success rate of the policy is then evaluated on the test set consisting of 1319 math questions, where for each question the success rate is evaluated using $50$ samples. We see a success rate amplification from $\pi_{\mathrm{ref}}$ originally (averaged on all prompts) at $21\%$ to $37.5\%$ at the end of the GRPO epoch.
Trajectory of Success rates Along GRPO Iterations
We randomly select few prompts from GSM8K test set and plot in Figure 5(a) the trajectory of the success rate of the model along the GRPO iteration (estimated from 50 samples from the model for each prompt). The success rate is computed from checkpoints of the model along the GRPO training. We see that the trajectory of the success rate $p(q)$ resembles the trajectory of a fixed point algorithm (see Figure 4 in Appendix LABEL:app:plots ). For some points the convegence is fast to the limit point $p^{*}=1$, for others we see an oscillatory behavior (similar to the one in last row in Figure 4). Interestingly when $p_{\mathrm{ref}}=0$, the probability of success does not move much along GRPO iterations as predicted by our theory.
(a) Success rate trajectory of the model on randomly selected prompts along the GRPO iteration.
(b) Convergence to fixed points
Figure 5. Fixed Points like behavior of success rate along GRPO iterations.
Convergence to Fixed Points of $h_{\varepsilon,p_{\mathrm{ref}}}$
We compute the success rate $p_{\mathrm{ref}}(q)$ for all prompts $q$ in the test of GSM8K, and compute the fixed point of $h_{\varepsilon,p_{\mathrm{ref}(q)}}(p)$ for all $q$ using a fixed point algorithm. This gives us $p^{*}(q)$ for all points in the test set. We compute the $\ell_{1}$ error between the success rate along the GRPO training $p_{n}(q)$ for a given prompt and the corresponding fixed point and plot the expected error on the prompts $\mathbb{E}_{q}|p_{n}(q)-p^{*}(q)|$ in Figure 5(b). We see that we have improvement in the convergence to fixed points up to a barrier due to computational and statistical errors as dicussed in Section 5. (Code to reproduce plots is in Supplementary material)
7. Conclusion
In conclusion, we have shown that GRPO with verifiable rewards can be viewed as an adaptive weighted contrastive loss (Section 2). We derived a closed-form recursion for the optimal policy, expressed in terms of the reference and old policies, and the probability of success (Section 3). This leads to a fixed-point equation, with GRPO’s probability of success converging to a fixed-point solution under mild assumptions (Section 4). Moreover, we proved that GRPO amplifies the probability of success compared to the reference model (Section 4). Finally, we showed that for approximate policies, the probability of success remains close to the fixed-point value as long as approximation statistical and optimization errors are small (Section 5).
Acknowledgement
We thank Omar El Mansouri for pointing out that no condition on $\beta$ is needed in Theorem 3.
Appendix A Broader Impact and Limitations
Our paper shows the GRPO advantage while simple enjoys a lot of good theoretical properties that ensure success rate amplification when using verifiable rewards. The main limitation in our work is that we study the non-clipped objective. This simplification makes the analysis possible and was used recently in [Vojnovic and Yun, 2025]. Extending the results to clipped objectives or to proximal objectives with $\mathsf{Kl}$ regularization with respect to previous policy remains challenging technically and we hope the techniques introduced in this work will help overcoming those technical challenges. The core contribution being theoretical we don’t see negative societal impact of our work.
Appendix B Assets
Hardware setup
All our experiments were run on one compute node with Dual 48-core Intel Xeon 8468, 2TB of RAM, 8 NVIDIA HGX H100 80GB SMX5, 8x 3.4TB Enterprise NVMe U.2 Gen4, and 10x NVIDIA Mellanox Infiniband Single port NDR adapters, running RedHat Enterprise Linux 9.5.
GRPO Config Setup
Libraries
Appendix C Proofs of Section 3
@@
[…]
@@
[…]
@@
Replacing the weights expressions from Equations (4) we obtain:
@@
We assume here that $10<p_{\mathrm{ref}}<$ 1.We can simplify $h_{\varepsilon}(p)$ as follows:
@@
Appendix D Proofs of Section 4

### v4

@@ 2.3. Stabilized GRPO
We will focus first on this non-clipped version.
2.4. GRPO Iterations
Algorithm 1 in Appendix B summarizes GRPO iterations (Stabilized and no clipping). We see that GRPO iterations can be written as a sequence of optimization resulting in policies we denote by $\pi_{\theta_{n}}$ the policy at iteration $n$. We see that GRPO iterations can be written for $n\geq 1$:
Note that in Algorithm 1, expectations are estimated using importance sampling from $\pi_{\theta_{n-1}}$, and each maximization problem is solved via gradient for $\mu$ steps. PoS are estimated using a group size $G$, i.e $G$ Monte-Carlo rollouts from $\pi_{\theta_{\text{old}}}(.|q)$.
$p_{n-1}(q)=\mathbb{E}_{o\sim\pi_{n-1}(.|q)}\mathbbm{1}_{r(q,o)=1}$ (7)
and the weights $\omega^{+}_{\varepsilon}$ and $\omega^{-}_{\varepsilon}$ are given in (4). We assume all throughout the paper that $\pi_{0}=\pi_{\mathrm{ref}}.$ Note that moving the optimization from a parametric space to the probability space can be seen as assuming that the hypothesis class of the parametric policies is large enough to represent all policies. Note that in GRPO iterations the policy at iteration $n$ depends upon the policy $\pi_{n-1}$ via the probability of success $p_{n-1}$, as well on the reference policy via the $\mathsf{KL}$ regularizer.
3. GRPO Dynamics: Fixed Point iteration for Probability of Success
Our goal in this Section is to analyze the dynamics of the GRPO iterations given in (GRPO Iterations).
Theorem 1 (GRPO Policy Dynamics).
Optimal GRPO iterations policies solving (GRPO Iterations) satisfy the following recursion, for $n\geq 1$:
where $Z_{n-1}(q)=p_{\mathrm{ref}}(q)\exp\left(\frac{1}{\beta}\omega^{+}_{\varepsilon}(p_{n-1}(q))\right)+(1-p_{\mathrm{ref}}(q))\exp\left(-\frac{1}{\beta}\omega^{-}_{\varepsilon}(p_{n-1}(q))\right),$ where the weights $\omega^{+}_{\varepsilon}$ and $\omega^{-}_{\varepsilon}$ are given in (13), the probability of success $p_{n-1}(q)$ of policy $\pi_{n-1}(\cdot|q)$ is given in (7), and $p_{\mathrm{ref}}(q)$ is the probability of success of the reference policy $\pi_{\mathrm{ref}}(\cdot|q)$: $p_{\mathrm{ref}}(q)=\mathbb{E}_{o\sim\pi_{\mathrm{ref}}(\cdot|q)}\mathbbm{1}_{r(q,o)=1}$.
Theorem 2 (GRPO’s Probability of Success Fixed Point Iteration).
Assume $p_{\mathrm{ref}}>0$, define for $\beta>0$:
$p_{n}(q)=h_{\varepsilon,p_{\mathrm{ref}}(q)}(p_{n-1}(q)),$ (8)
and $p_{0}(q)=p_{\mathrm{ref}}(q)$.
Remark 1 (Importance of $\varepsilon>0$).
Note if $\varepsilon=0$, $h_{\varepsilon,p_{\mathrm{ref}}}$ is no longer continuous at 0 and $1$ and we can no longer guarantee existence of fixed points on $[0,1]$.
We study in the following proposition properties of the function $h_{\varepsilon,p_{\mathrm{ref}}}$:
Proposition 1 (Properties of $h_{\varepsilon,p_{\mathrm{ref}}}$).
$h_{\varepsilon,p_{\mathrm{ref}}}$ satisfies the following properties:
  • Existence of fixed points: $h_{\varepsilon,p_{\mathrm{ref}}}$ is continuous $from[0,1]$ to $[0,1]$ and hence admits at least a fixed point $p^{*}$ in $[0,1]$ (no guarantees for a unique fixed point)
  • Let $\operatorname{logit}(p)=\log\!\left(\frac{p}{1-p}\right),\sigma(x)=\frac{1}{1+e^{-x}},((\operatorname{logit}\circ\sigma)(x)=x,\ (\sigma\circ\operatorname{logit})(p)=p).$ Define $\Omega_{\varepsilon}(p)=\omega^{+}_{\varepsilon}(p)+\omega^{-}_{\varepsilon}(p)=(p(1-p)+\varepsilon)^{-\frac{1}{2}}.$ We have:
    $h_{\varepsilon,p_{\mathrm{ref}}}(p)=\sigma\!\left(\operatorname{logit}\big(p_{\mathrm{ref}}\big)+\frac{\Omega_{\varepsilon}(p)}{\beta}\right).$
We drop in the sequel $q$, when referring to the sequence $p_{n}(q)$, and write for short $p_{n}$. If the sequence defined in GRPO’s probability of success iteration (8) converges we have therefore by continuity of $h_{\varepsilon,p_{\mathrm{ref}}}$:
3.1. GRPO: Fixed Point Iteration and Success Amplification
Note that from the third item in proposition 1 the PoS recurrence in Theorem 2 can be written in terms of success odds as follows:
$\operatorname{logit}\big(p_{n}(q)\big)=\operatorname{logit}\big(p_{\mathrm{ref}}(q)\big)+\frac{\Omega_{\varepsilon}(p_{n-1}(q))}{\beta}$
Theorem 3 (GRPO amplifies the probability of success).
For $q\sim\rho_{\mathcal{Q}}$ assume $0<p_{\mathrm{ref}}(q)<1$. Let $p^{*}(q)$ be a fixed point of $h_{\varepsilon,p_{\mathrm{ref}}(q)}$ we have $p^{*}(q)>p_{\mathrm{ref}}(q)$.
We see from Theorem 3 for any prompt $q$, the fixed point PoS $p^{*}(q)$ of the GRPO iteration leads to an amplification of the probability of success of the reference model $p_{\mathrm{ref}}(q)$. Note if $p_{\mathrm{ref}}(q)=0$ or $p_{\mathrm{ref}}(q)=1$ , the iteration will lead to $p^{*}(q)=0$ and $p^{*}(q)=1$ respectively. In this case the fixed point is not necessarily stable and a condition on $\beta$ is needed for its stability (See appendix C.2 )
4. Mirror GRPO: Mirror Descent with GRPO Calibrated Reward
Note that we previously considered GRPO with no-clipping and with a $\mathsf{KL}$ regularization to $\pi_{\mathrm{ref}}$.
We consider here a mirror GRPO with a regularization to $\pi_{n-1}$ in addition to $\pi_{\mathrm{ref}}$. For $n\geq 1$:
$\max_{\pi}\;\mathbb{E}_{q\sim\rho_{\mathcal{Q}}}\left(\mathbb{E}_{\pi(\cdot|q)}A_{n-1}(q,\cdot)-\;\beta\,\left(\alpha\mathsf{KL}\!\Big(\pi(\cdot\mid q)\,\Big\|\,\pi_{\mathrm{ref}}(\cdot\mid q)\Big)+(1-\alpha)\mathsf{KL}\!\Big(\pi(\cdot\mid q)\,\Big\|\,\pi_{n-1}(\cdot\mid q)\Big)\right)\right),$ (9)
where:
$A_{n-1}(q,o)=\begin{cases}+\omega^{+}_{\varepsilon}(p_{n-1}(q)),&r(q,o)=1\\[2.0pt] -\omega^{-}_{\varepsilon}(p_{n-1}(q)),&r(q,o)=0,\end{cases}$ (10)
and $p_{n-1}(q)=\mathbb{P}_{\pi_{n-1}(\cdot|q)}(r(q,o)=1)$, and $\pi_{0}=\pi_{\mathrm{ref}}$.
If $\alpha=1$ we obtain $\mathsf{KL}$ regularization to the $\pi_{\mathrm{ref}}$. If $\alpha=0$, we obtain mirror regularization to the previous iteration without considering the reference. Many recent works suggested using $\alpha=0$ such as DAPO (Yu et al., 2025) i.e removing the regularization to the reference in GRPO while maintaining the clipping. Note that proximal methods with regularization to previous iterates play the same role of clipping (Tomar et al., 2021; Gunter et al., 2024). Indeed PPO style clipping (Schulman et al., 2017) was introduced as an approximation of such proximal mirror descent.
We study in the following the case $\alpha=0$, the general case $\alpha>0$ is analyzed in Appendix E.
Theorem 4 gives the optimal policy for Mirror-GRPO iterations, and its corresponding PoS recurrence:
Theorem 4 (Mirror-GRPO, $\alpha=0$).
Fix $\alpha=0$ and a prompt $q$ and let $\beta>0$. Let $\Omega_{\varepsilon}(p)=\frac{1}{\sqrt{p(1-p)+\varepsilon}}.$ Then the following holds:
  1. (1)
    Optimal policy. The maximizer of (9) is
    $\pi_{n}(o|q)=\frac{1}{Z_{n-1}(q)}\pi_{n-1}(o|q)\exp\left(\frac{1}{\beta}\left(\omega^{+}_{\varepsilon}(p_{n-1}(q))\mathbbm{1}_{r(q,o)=1}-\omega^{-}_{\varepsilon}(p_{n-1}(q))\mathbbm{1}_{r(q,o)=0}\right)\right),$
    where $Z_{n-1}(q)=p_{n-1}(q)\exp\left(\frac{1}{\beta}\omega^{+}_{\varepsilon}(p_{n-1}(q))\right)+(1-p_{n-1}(q))\exp\left(-\frac{1}{\beta}\omega^{-}_{\varepsilon}(p_{n-1}(q))\right).$
  2. (2)
    PoS and odds recurrences. The PoS of $\pi_{n}(\cdot|q)$, $p_{n}(q)$, satisfies the following recurrence:
    $\operatorname{logit}(p_{n}(q))$ $=\operatorname{logit}(p_{n-1}(q))\;+\;\frac{\Omega_{\varepsilon}(p_{n-1}(q))}{\beta},$ (11)
    $p_{n}(q)$ $=h_{\varepsilon,\beta}(p_{n-1}(q))=\sigma\!\Big(\operatorname{logit}(p_{n-1}(q))+\Omega_{\varepsilon}(p_{n-1}(q))/\beta\Big).$ (12)
When compared with Theorem 2, we see that $p_{n-1}(q)$, replaces $p_{\mathrm{ref}}(q)$ in the logit inside the sigmoid.
Theorem 5 (Monotone Improvement and Absorbing Fixed Points).
Fix a prompt $q$,the PoS iterations $p_{n}(q)$ of Mirror-GRPO ($\alpha=0$) have the following properties:
  1. (1)
    Monotone improvement and absence of interior fixed points. For any $p_{n-1}\in(0,1)$, $\Omega_{\varepsilon}(p_{n-1})/\beta>0$, hence $\operatorname{logit}(p_{n})>\operatorname{logit}(p_{n-1})$ and $p_{n}>p_{n-1}$. Consequently, the equation $p=\sigma(\operatorname{logit}(p)+\Omega_{\varepsilon}(p)/\beta)$ has no solution in $(0,1)$. The only fixed points are at the boundary: $p\in\{0,1\}$.
  2. (2)
    Convergence and stability. The fixed points of Mirror-GRPO iterations ($\alpha=0$) satisfy:
    1. (a)
      If $p_{\mathrm{ref}}(q)\in(0,1)$, then $(p_{n}(q))_{n}$ is strictly increasing and bounded by $1$, hence $p_{n}\uparrow 1$.
    2. (b)
      If $p_{\mathrm{ref}}(q)\in(0,1)$, $p^{*}=1$ is (globally) stable fixed point: $\lim_{n\to\infty}p_{n}(q)=1$.
    3. (c)
      If $p_{\mathrm{ref}}(q)=0$ then $p_{n}(q)=0$ for all $n$.
When compared with Theorem 3, we see for non-zero $p_{\mathrm{ref}}(q)$, Mirror-GRPO iterations of probability of success converges to $1$ that is a stable fixed point, whereas for GRPO with only reference regularization we may have an interior fixed point $p^{*}(q)>p_{\mathrm{ref}}(q)$. In both case for zero $p_{\mathrm{ref}}(q)$, GRPO with reference regularization or Mirror GRPO don’t create successes, and the fixed point success remains at zero. From a practical point of view removing the reference regularization is convenient as one does not need to keep in memory the reference model in addition to the current model . In addition it has more favorable PoS guarantees than reference regularization only. Nevertheless in many situations one wants to achieves good performance on a task via RL training while maintaining the knowledge of the reference model and hence the case $\alpha>0$ is also of interest, we study this case fully in Appendix E. The main takeaway in that scenario where interpolate between $\alpha=0$ and $\alpha=1$, is that we lose monotonic improvement. The PoS iteration incurs what we call a Rényi correction that encodes the mismatches in success and failures between the reference and the previous iteration and we are back to an interior fixed point in $(0,1)$ and no guarantees of global stability as in the mirror-GRPO case.
5. Dr. GRPO and mean-only Normalization
We turn now to another reward normalization proposed in Dr. GRPO (Liu et al., 2025). Liu et al. (2025) suggests to use a mean-only normalization in GRPO. In our notations this corresponds to the following reward calibration:
$A_{\pi_{\theta_{\text{old}}}}(q,o)=\begin{cases}+\omega^{+}(p(q)),&r(q,o)=1,\\ -\omega^{-}(p(q)),&r(q,o)=0,\end{cases}\quad\omega^{+}(p)=1-p,\qquad\omega^{-}(p)=p.$ (13)
This results in the following (no clipping) Dr. GRPO iterations for PoS:
$\operatorname{logit}(p_{n}(q))=\operatorname{logit}(p_{\mathrm{ref}}(q))+\frac{1}{\beta}.$
and the following for Mirror Dr. GRPO
$\operatorname{logit}(p_{n}(q))=\operatorname{logit}(p_{n-1}(q))+\frac{1}{\beta}.$
These expressions can be obtained applying Theorem 8 in Appendix D for this particular weighting with $\Omega(p)=1$.
When Compared with (no) clip GRPO, DR. GRPO has a trivial constant fixed point $p^{*}(q)=\sigma(\operatorname{logit}p_{\mathrm{ref}}(q)+\frac{1}{\beta})$. While for Mirror Dr. GRPO , $L_{n}(q)=\operatorname{logit}(p_{n}(q))$ is an arithmetic progression and $L_{n}(q)=L_{\mathrm{ref}}(q)+\frac{n}{\beta}$ and $p_{n}(q)\uparrow 1$ for non degenerate $p_{\mathrm{ref}}(q)\in(0,1)$. Comparing to Mirror GRPO we have a similar convergence to a PoS of 1 but the iteration are adaptive in the case of Mirror GRPO:
$\operatorname{logit}(p_{n}(q))=\operatorname{logit}(p_{n-1}(q))+\frac{\Omega_{\varepsilon}(p_{n-1}(q))}{\beta}=\operatorname{logit}(p_{n-1}(q))+\frac{1}{\beta(\sigma^{2}_{n-1}(q)+\varepsilon)^{\frac{1}{2}}},$
we can think that the variance normalization corresponds to mean-only normalization with an adpative effective $\beta_{\text{eff}}=\beta\sqrt{\sigma^{2}_{n-1}(q)+\varepsilon}$. For low variance we make large increments in the logits of PoS and for high variance, we make smaller increments in the logits of PoS.
6. Discussion and Conclusion
Table 1 in the Appendix summarizes different flavors of GRPO we studied in this paper and their corresponding probability of success iterations.
The main dimensions these variants differ on are: 1) the reward calibration: mean and variance normalization as in the original GRPO or mean-only normalization as in Dr GRPO (Liu et al., 2025). Our theory showed that the normalization results in different weighting schemes, non linear in the PoS for GRPO and linear in the PoS for Dr GRPO. 2) As discussed earlier the analysis of the PPO style clipping to maintain the policy updates in the vicinity of the old policy is challenging and it has been shown to be more stable to use mirror policy descent to train LLMs with RL (Gunter et al., 2024). Hence we distinguish GRPO variants also with respect to the anchor distribution on which the KL regularization is applied : no-clip refers to $\pi_{\mathrm{ref}}$ regularization only. Mirror corresponds to the KL regularization given in (9) with respect to the previous iterate ($\alpha=0$), we also consider the regularization to both reference and previous iteration (two-KL) for $\alpha>0$. For $\alpha=0$, we see that we obtain a monotonic improvement in the PoS whereas mixing the reference and the previous iterate in the iterations does not guarantee monotonic improvement. The PoS iteration in this case depends on the mismatch in success and failures between the reference and the previous iteration that we quantify in Appendix E via a Rényi correction.
From a practical point of view Table 1 suggests the following in using GRPO in training LLMs:
Practical Takeaways Normalization equivalence. The mean+variance normalization in GRPO is equivalent from PoS point of view to mean-only normalization using an adaptive KL regularization $\beta_{\mathrm{eff}}=\beta\sigma(q)$. One can use either a fixed $\beta$ and get constant increments in log PoS odds via mean-only calibration, or use mean calibration with $\beta_{\mathrm{eff}}$ as a KL regularizer which results in adaptive increments that are equivalent to mean + variance normalization without having to divide by the variance in the advantage. Mirror versus Clipping and Reference Mixing Mirror GRPO (KL to previous iteration only ) instead of clipped GRPO guarantees monotonic improvement and convergence to PoS of 1 for non degenerate $p_{\mathrm{ref}}$. Mirror GRPO has the best theoretical and practical guarantees. Adding the reference regularization to this mirror descent results in an internal fixed point and no monotonic improvement is guaranteed. Practically speaking, keeping a reference policy in memory increases bandwidth/latency and can slow training for large models. Coverage and exploration. In all cases GRPO does not create successes and 0 is an absorbing fixed point if $p_{\mathrm{ref}}(q)=0$. Hence it is important to maintain successes exploration (e.g., temperature, entropy bonus, or data mixing) so successes have nonzero support.
Appendix A Experimental Validation
Setup We use the GSM8K dataset from Cobbe et al. [2021] (MIT license), and
Qwen/Qwen2.5-0.5B-Instruct (Apache 2.0 license) by Yang et al. [2024] as the reference policy. We use GRPO implementation in TRL [von Werra et al., 2020b], and train on the training split of GSM8K on a node with 8 GPUs (GPU0 for the vLLM server and 7 other GPUs for distributed training). We use a learning rate $5e^{-6}$, clipping $\varepsilon=0.2$ and the KL regularizer $\beta=0.1$, and $\mu$ in Algorithm 1 is set to $\mu=10$. Other hyperparameters are given in Appendix H . We use the correctness of the LLM output as a reward.
Success Rate Amplification The success rate of the policy is then evaluated on the test set consisting of 1319 math questions, where for each question the success rate is evaluated using $50$ samples. We see a success rate amplification from $\pi_{\mathrm{ref}}$ originally (averaged on all prompts) at $21\%$ to $37.5\%$ at the end of the GRPO epoch.
Trajectory of Success rates Along GRPO Iterations We randomly select few prompts from GSM8K test set and plot in Figure 3 the trajectory of the success rate of the model along the GRPO iteration (estimated from 50 samples from the model for each prompt). The success rate is computed from checkpoints of the model along the GRPO training. We see that the trajectory of the success rate $p(q)$ resembles the trajectory of a fixed point algorithm (see Figure 5 in Appendix G ). For some points the convergence is fast to the limit point $p^{*}=1$, for others we see an oscillatory behavior (similar to the one in last row in Figure 5). Interestingly when $p_{\mathrm{ref}}=0$, the probability of success does not move much along GRPO iterations as predicted by our theory.
Figure 3. Success rate trajectory of the model on randomly selected prompts along the GRPO iters.
Appendix B Algorithm
@@
Algorithm 1 Iterative GRPO with verifiable rewards, modified from [Shao et al., 2024]
@@
8:  Compute $\hat{A}(q,o_{i})$ using (4), where $\hat{p}(q)=\hat{p}_{\theta_{\text{old}}}(q)=\frac{1}{G}\sum_{i=1}^{G}\mathbbm{1}_{r(q,o_{i})=1}$
@@
10:    Update the policy model $\pi_{\theta}$ by maximizing GRPO objective with gradient ascent
11:  end for
12:end for
13:Output $\pi_{\theta}$
Appendix C (No-Clipping ) GRPO : Proofs of Section 3
@@
$Z_{n-1}(q)$ $=\int d\pi_{\mathrm{ref}}(o|q)\exp\left(\frac{1}{\beta}\left(\omega^{+}_{\varepsilon}(p_{n-1}(q))\mathbbm{1}_{r(q,o)=1}-\omega^{-}_{\varepsilon}(p_{n-1}(q))\mathbbm{1}_{r(q,o)=0}\right)\right)$
@@
$=\frac{1}{Z_{n-1}(q)}\int d\pi_{\mathrm{ref}}(o|q)\exp\left(\frac{1}{\beta}\left(\omega^{+}_{\varepsilon}(p_{n-1}(q))\mathbbm{1}_{r(q,o)=1}-\omega^{-}_{\varepsilon}(p_{n-1}(q))\mathbbm{1}_{r(q,o)=0}\right)\right)\mathbbm{1}_{r(q,o)=1}$
@@
Replacing the weights expressions from (13) we obtain:
@@
We assume here that $0<p_{\mathrm{ref}}<1$.We can simplify $h_{\varepsilon}(p)$ as follows:
@@
Turning to third point:
$h_{\varepsilon,p_{\mathrm{ref}}}(p)$ $=\sigma\left(z(p)\right)$
$=\sigma\left(\log\frac{p_{\mathrm{ref}}}{1-p_{\mathrm{ref}}}+\frac{1}{\beta}\frac{1}{\sqrt{p(1-p)+\varepsilon}}\right)$
$=\sigma\left(\operatorname{logit}(p_{\mathrm{ref}})+\frac{1}{\beta}\frac{1}{\sqrt{p(1-p)+\varepsilon}}\right)$
and hence:
$\operatorname{logit}\left(h_{\varepsilon,p_{\mathrm{ref}}}(p)\right)=\operatorname{logit}(p_{\mathrm{ref}})+\frac{1}{\beta}\frac{1}{\sqrt{p(1-p)+\varepsilon}}.$ (15)
C.1. Proofs of Section 3.1
@@
C.2. Stability fixed Point GRPO with Reference Only Regularization
We drop in the sequel the dependency on $q$ to simplify notations and turn to the second question regarding the convergence of the GRPO sequence given in (8) to a fixed point $p^{*}$ of $h_{\varepsilon,p_{\mathrm{ref}}}$. Given the properties of $h_{\varepsilon,p_{\mathrm{ref}}}$, we can characterize the limit point of the GRPO iteration as $n\to\infty$ as follows, as a consequence of the local Banach fixed-point theorem:
Theorem 6 (Local Fixed Point Convergence).
$\lim_{n\to\infty}p_{n}=p^{*}.$
Lemma 1.
$=p^{*}(1-p^{*})\frac{2p^{*}-1}{2\beta\,[p^{*}(1-p^{*})+\varepsilon]^{3/2}}\$
We see from Figure 4 in Appendix G the lower bound on $\beta$ required to ensure local convergence of GRPO iterations to a fixed point $p^{*}$. Figure 5 in Appendix G shows iteration (8) as a function of $n$ for different values of $\beta$ and $p_{\mathrm{ref}}$. We see that in most cases, there is a sharp transition where we observe fast convergence to $1$ or to a fixed point $p^{*}$. For $\beta=5$ and $p_{\mathrm{ref}}=0.001$, we see a divergent behavior.
Remark 2.
Practical Implications.
In practical implementations GRPO is applied successively in stages where $\pi_{\mathrm{ref}}$ is set to the last iteration from the GRPO training in each stage [Shao et al., 2024]. In light of our theory this ensures that we are amplifying the probability of success w.r.t the new $\pi_{\mathrm{ref}}$, coming the previous GRPO stage.